OpenStudy (anonymous):

Michael and Kathryn bowl together and their combined total score for one game was 425 points. Michael’s score was 70 less than twice Kathryn’s. What were their scores? Which is a system of equations to model the problem if x represents Kathryn’s score and y represents Michael’s score? (Points : 1)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

x + y = 70 y = 2x – 425 x – y = 425 y = 2x – 70 x + y = 425 y = 2x – 70 x + y = 425 y = 2x + 70 these are the options

OpenStudy (anonymous):

let x=Kathryn and y=Micheal 425=x+y y=2x-70

OpenStudy (anonymous):

x + y = 425 y = 2x - 70 I agree with sodukogirl :)

OpenStudy (anonymous):

thank you